OpenStudy (calculator):

Integrating problem x/(1-x)^2

OpenStudy (calculator):

Wait me show my steps

OpenStudy (anonymous):

u sub right ?

OpenStudy (calculator):

\[x=\tan \theta\]\[dx=\sec ^{2}\theta d \theta\] \[\int\limits_{}^{}\frac{\tan \theta}{(1-\tan \theta)^{2}}\sec ^{2}\theta d \theta\] \[\int\limits\limits\limits_{}^{}\frac{\tan \theta \sec ^{2}\theta}{(1+\tan ^{2} \theta-2\tan \theta)}\ d \theta\] \[\int\limits\limits\limits\limits_{}^{}\frac{\tan \theta \sec ^{2}\theta}{(\sec ^{2}\theta-2\tan \theta)}\ d \theta\] Im stucked here

OpenStudy (anonymous):

try \[u=1-x\] first

OpenStudy (agreene):

have you learned partial fractions yet? thats the method I used on this.

OpenStudy (agreene):

with u=1-x ;)

OpenStudy (agreene):

er x-1 rather.

OpenStudy (anonymous):

\[u=1-x, x=1-u\] \[-du=dx\]

OpenStudy (anonymous):

you get \[-\int \frac{1-u}{u^2}du\]

OpenStudy (anonymous):

you can also do it by partial fractions, and also by a gimmick

OpenStudy (calculator):

Oh okay, i get the answer, but the problem is with myself , i dont know when to use the trigonometry :(

OpenStudy (agreene):

\[\int \frac{x}{(x-1)^2} dx = \int (\frac{1}{x-1}+\frac1{(x-1)^2})dx\] let: u = x-1 du=dx plug and chug.

OpenStudy (anonymous):

@green that would be probably the most work

OpenStudy (agreene):

probably, but i like partial fractions :)

OpenStudy (anonymous):

more than one way to skin a cat

OpenStudy (anonymous):

plus you skipped a step !

OpenStudy (calculator):

i dont know when to use the trigonometry :(

OpenStudy (agreene):

@Calculator Once you know all the methods of integration, it takes some time to get to know when to use what--and even then, people argue about when to do what... kinda like me and @satellite73 ... fairly often, lol

OpenStudy (anonymous):

how do you know \[\frac{x}{(1-x)^2}=\frac{1}{1-x}+\frac{1}{(1-x)^2}\] and is it even true??

OpenStudy (agreene):

partial fractions definition QED

OpenStudy (anonymous):

to be honest you usually use a trig sub when you come to the section in the book on trig subs

OpenStudy (calculator):

but the thing is in the exam they didnt tell you use which technique, lol

OpenStudy (anonymous):

then you will see some similar forms after that

OpenStudy (agreene):

True, but at a certain point, you need to use all of the methods... if you have a worthwhile calc II teacher they will construct problems that utilize all the methods in a single question to truly test you.

OpenStudy (anonymous):

yeah nothing more annoying that techniques of integration bunch of teachers showing off look we can do this techninque look we can do that one truth is almost all functions are not integrable

OpenStudy (agreene):

lol i dont even use integration anymore, functions that are continuous and easily defined are so rare in actual applications.

OpenStudy (callisto):

i don't know if it works, you can use trigo sub when you see the form \[\sqrt{a ^{2}-x ^{2}}\]or\[\sqrt{x ^{2}-a ^{2}}\] or \[\sqrt{a ^{2}+x ^{2}}\] (ps i'm really bad in dealing with integration)